# TriVLA: A Triple-System-Based Unified Vision-Language-Action Model for General Robot Control

### v2

@@ 5.1 Simulation Setups and Baselines
Comparison methods. Generalist robot policies have been extensively investigated in prior research. In our experiments, we select a representative subset of prior methods for comparison, focusing on those that have achieved state-of-the-art performance or employ approaches similar to ours.
@@
Figure 4: Qualitative case study. Our TriVLA performs exceptionally well in long-horizon missions. Taking the CALVIN simulation task as an example, TriVLA integrates world knowledge for intent understanding and leverages a world model for future prediction when given multiple sequential instructions, enabling effective execution of long-horizon tasks.
Quantitative Results. Comparisons on the Calvin benchmark are presented in Table 1. Experimental results for Robo-Flamingo, Susie, GR-1, and 3D Diffuser Actors are extracted from their respective original publications. MDT results are obtained from the official implementation. RT-1 and UniPi results are sourced from Li et al. (2023b) and Black et al. (2023), respectively. We additionally executed the Diffusion Policy using the official open-source codebase with CLIP language conditioning. Our proposed TriVLA significantly surpasses previous state-of-the-art results. Remarkably, despite training on only 10% of the annotated Calvin ABC dataset, our method achieved an average task completion length of 3.46, outperforming related methods trained on the full dataset. Furthermore, TriVLA attained the highest performance on the MetaWorld benchmark comprising 60 tasks, as detailed in Table 2. It outperforms the strongest VPP baseline in average success rate. Quantitative results on the LIBERO benchmark are presented in Table 3, where each method is evaluated over 500 trials per task suite using 3 random seeds. The results demonstrate that TriVLA effectively adapts to LIBERO simulation tasks, achieving best or competitive performance relative to baseline methods.
Visualization of trajectories. We provide three qualitative examples of action sequences generated by TriVLA in Figure 4. Given multiple consecutive instructions—for instance, “Pull the handle to open the drawer,” “Grasp and lift the pink block,” “Use the switch to turn on the light bulb,” and “Store the grasped block in the sliding cabin”—TriVLA demonstrates the ability to comprehend instructions, infer intent, and utilize predictive capabilities to accomplish long-horizon tasks. The results demonstrate that TriVLA employs VLMs and VDMs for both high-level reasoning based on common knowledge and dynamic predictive representation provided by a world model. TriVLA integrates world knowledge to enhance intent understanding and utilizes a world model for future state prediction when processing multiple sequential instructions, thereby enabling effective long-horizon task execution.
@@
Figure 5: Visualization of one-step visual representations of dynamics perception module. We can observe that representation can provide valuable information on physical dynamics, although the textures and details are not precise enough.
5.2 Ablation Study
@@
Visualization of Dynamics Perception. We employ a stable video diffusion model as the dynamics perception module, executing a single forward pass to obtain visual representations that encompass both current static information and predicted future dynamics. As illustrated in Figure 5, we present visualizations of ground-truth futures alongside single-step predictions on the Bridge benchmark Walke et al. (2023). The visualization results indicate that single-step representations convey critical information, including object and robot arm motion, thereby effectively supporting downstream action learning. The dynamics perception module models entire video sequences and predicts future frames conditioned on current observations and instructions, demonstrating a sound understanding of physical dynamics.
Effectiveness of Vision-Language Module. The System 2 vision-language module (VL-M) of TriVLA is a pre-trained Vision-Language Model (VLM) responsible for processing the robot’s visual perception and language instructions to interpret the environment and comprehend task goals. Experiments were conducted to evaluate the effectiveness of System 2. As summarized in Table 4, integration of the vision-language module improved performance from 4.06 to 4.37, with inference time increasing from 136 ms to 155 ms. These results suggest that integrating System 2 significantly enhances the accuracy of action generation.
Effectiveness of Dynamics Perception Module. The System 3 dynamics perception module (DP-M) is a general-purpose video diffusion model fine-tuned on internet-sourced human and robot manipulation datasets. This module is designed to develop a controllable video generation model enhancing predictive capabilities within the manipulation domain. To assess its effectiveness, we perform an ablation study by integrating the dynamics perception module into the proposed TriVLA framework. The results, summarized in Table 4, were obtained using a single NVIDIA H100 GPU. Notably, incorporation of the dynamics perception module results in a significant enhancement of robot control performance.
Results on Calvin Benchmark
Sub-system Performance
VL-M DP-M PL-M Avg. Length $\uparrow$ Latency $\downarrow$ Params $\downarrow$
3.68 29.29ms 0.53B
4.06 115.19ms 1.87B
4.37 142.69ms 3.39B
Table 4: Sub-system ablation studies. We report the average success length and latency on the CALVIN simulation benchmark. VL-M denotes the System 2 vision-language module, DP-M denotes the System 3 dynamics perception module, and PL-M denotes the System 1 policy learning module.
6 Conclusion
TriVLA presents a novel triple-system architecture that effectively integrates vision-language understanding with dynamic perception, enhancing the robot’s capability to capture both static and future dynamic information. This integration enables more fluid and generalized control policies for robot manipulation. Experimental results demonstrate that TriVLA outperforms state-of-the-art imitation learning baselines on standard simulation benchmarks and challenging real-world tasks, highlighting its potential for broad applications in general robotic control.
7 Acknowledgments
This work was supported by the Science and Technology Commission of Shanghai Municipality (No. 24511103100) and in part by NSFC Project (62176061), and Shanghai Technology Development and Entrepreneurship Platform for Neuromorphic and AI SoC.

### v3

@@ Appendix D Comparison Methods
Generalist robot policies have been extensively investigated in prior research. In our experiments, we select a representative subset of prior methods for comparison, focusing on those that have achieved state-of-the-art performance or employ approaches similar to ours.
@@
Figure 7: Visualization of Episodic Dynamics Perception on the Bridge Benchmark. The red box indicates one-step prediction, the blue box corresponds to full-step prediction, and the green box marks the ground truth. We can observe that the representation can provide valuable information on physical dynamics, although the textures and details are not precise enough.
Appendix E Details and More Results of Episodic Dynamics Perception
@@
We employ a stable video diffusion model as the dynamics perception module, performing a single forward pass to obtain visual representations that encompass both current static information and predicted future dynamics. As illustrated in Figure 7, we present visualizations of ground-truth futures alongside single-step and full-step predictions on the Bridge benchmark. The visualization results indicate that single-step representations convey critical information, including object and robot arm motion, thereby effectively supporting downstream action learning. The dynamics perception module models entire video sequences and predicts future frames conditioned on current observations and instructions, demonstrating a sound understanding of physical dynamics.
Figure 8: Visualization of Episodic Dynamics Perception on Real-world Tasks. The blue box corresponds to the full-step prediction, and the green box marks the ground truth of the current timestep. We can observe that representation can provide valuable information on physical dynamics, although the textures and details are not precise enough.
Additionally, we extend our analysis to real-world experiments, where we replicate the same predictive framework on an actual robot. As shown in Figure 8 and  9, the true-to-life experimental results mirror the findings from the simulated setup, further validating the robustness of our prediction model. These real-world predictions are crucial, as they show that our system not only captures the motion of objects and the robot arm but also adapts to real-world uncertainties, such as sensor noise and minor mechanical inaccuracies.
Figure 9: Visualization of Episodic Dynamics Perception on Real-world Tasks. The red box indicates the one-step prediction, the blue box corresponds to the full-step prediction, and the green box marks the ground truth. We can observe that representation can provide valuable information on physical dynamics, although the textures and details are not precise enough.
The real-world prediction visualizations highlight several important aspects of our approach:
  • Accurate Motion Forecasting: Even in the face of real-world complexities, the system accurately predicts future motions of both the robot arm and surrounding objects. This is key for enabling high-level decision-making and adaptive action execution.
  • Real-World Generalization: The model demonstrates strong generalization capabilities, transferring learned predictions from the benchmark environment to practical settings without requiring extensive retraining. The system’s robustness in handling real-world dynamics proves the versatility of the proposed architecture.
By leveraging these visualizations and predictions in both the simulated and real environments, we show that our framework can bridge the gap between theoretical modeling and real-world robotic applications. This provides a powerful tool for task generalization, enabling robots to efficiently plan and execute complex actions in diverse scenarios. In summary, the combination of simulated and real-world results not only validates the robustness of our prediction framework but also underscores its potential for real-time action learning and autonomous decision-making in physical environments. The visualizations of future predictions further support the importance of incorporating dynamic modeling into robotic systems, fostering a deeper understanding of physical interactions and improving the overall system performance.
Figure 10: Qualitative case study of CALVIN benchmark. Our TriVLA performs strongly in long-horizon missions. For example, in the CALVIN simulation task, it integrates world knowledge to interpret intent and uses a world model to predict future states. Given multiple sequential instructions, TriVLA can effectively execute long-horizon tasks.
Appendix F Qualitative Analysis and Results
We provide qualitative examples of action sequences generated by TriVLA in Figure 10. Given multiple consecutive instructions—for instance, “Pull the handle to open the drawer,” “Grasp and lift the pink block,” “Use the switch to turn on the light bulb,” and “Store the grasped block in the sliding cabin”—TriVLA demonstrates the ability to comprehend instructions, infer intent, and utilize predictive capabilities to accomplish long-horizon tasks. The results demonstrate that TriVLA employs VLMs and VDMs for both high-level reasoning based on common knowledge and dynamic predictive representation provided by a world model. TriVLA integrates world knowledge to enhance intent understanding and utilizes a world model for future state prediction when processing multiple sequential instructions, thereby enabling effective long horizon task execution.
  • Precise Alignment with Modeled Dynamics: Since simulation provides deterministic or near-deterministic dynamics, TriVLA demonstrates highly consistent outcomes across repeated trials. For instance, during multi-step manipulation sequences, object trajectories match predicted states almost perfectly, showcasing the model’s capability to exploit stable environments for accurate planning.
  • Stress Testing under Controlled Perturbations: Simulation allows for the systematic injection of domain variations, such as randomized object masses, altered friction coefficients, or unexpected collisions. TriVLA adapts to these controlled perturbations by updating its predictions accordingly, highlighting its resilience under a wide spectrum of simulated uncertainties.
  • Long-Horizon Reasoning at Scale: Most importantly, simulation provides an ideal platform for TriVLA to validate long-horizon planning strategies across diverse scenarios. By anticipating future states over extended horizons, the model learns to optimize its policy in a wide variety of contexts, generating transferable skills that can later be fine-tuned for real-world execution.
Appendix G Real-world experiments
In addition to the simulated results, we also conducted real-world experiments, where TriVLA successfully generates and executes action sequences on a physical robot. As shown in Figure 11, the robot accurately follows the same sequence of tasks, starting from pulling the handle to opening the drawer, grasping and lifting the pink block, using the switch to turn on the light bulb, and finally storing the block in the sliding cabin. These real-world results closely align with the predictions and actions generated in the simulated environment, further validating TriVLA’s ability to handle complex, multi-step tasks in dynamic, real-world scenarios.
Figure 11: Qualitative case study of real-world tasks. Our TriVLA performs well in real-world tasks, successfully executing both short-horizon and long-horizon manipulations. The results illustrate its ability to integrate perception, prediction, and control for reliable task completion under real-world conditions.
The real-world experiments highlight several key advantages of TriVLA’s approach:
  • Real-Time Sequence Execution: The robot efficiently processes and executes long-horizon tasks in real-time, leveraging TriVLA’s ability to predict intermediate states and adjust actions accordingly. Despite the inherent variability and unpredictability of the real world, such as slight environmental changes or sensor noise, TriVLA’s predictive capabilities allow the robot to remain on task without requiring extensive retraining.
  • High Fidelity in Task Completion: As the robot progresses through the sequence of actions, it demonstrates a strong alignment between predicted outcomes and actual results. For instance, after pulling the handle, the drawer opens correctly, and the robot adjusts its grip on the block while maintaining stability during the lift. This showcases the robustness of TriVLA’s predictive world model in real-world settings.
  • Dynamic Adaptation to Uncertainty: The real-world setup also presents challenges like minor inaccuracies in motor control or shifting environmental conditions. TriVLA exhibits impressive adaptability, dynamically adjusting predictions and actions to account for these uncertainties, ensuring continued task success.
  • Long-Horizon Task Planning: Perhaps most notably, TriVLA demonstrates its ability to execute long-horizon plans by integrating both episodic memory and predictive reasoning. By leveraging its world model, TriVLA is able to anticipate future states and proactively adjust actions, ensuring that all steps of the sequence are successfully completed, even in the presence of unforeseen challenges.
Overall, these real-world experiments reinforce the core strength of TriVLA: its ability to understand complex instructions, reason about sequential actions, and predict future states—an essential capability for enabling embodied agents to perform sophisticated tasks autonomously and effectively in the real world. Through the combination of simulated and real-world action sequence generation, TriVLA proves to be a highly capable architecture for long-horizon task execution, paving the way for more advanced and adaptable autonomous systems.